# Graph Reconstruction from Noisy Random Subgraphs

### v1

@@ I Introduction
We are interested in the number of independently generated traces that are necessary to reconstruct the graph $G$ (with high probability). We refer to this number as the sample complexity of reconstruction. The problem was studied by McGregor and Sengupta [1], who considered the noiseless setting where $p_{e}=1$ (or equivalently $f_{e}=0$). They showed that $O(p_{v}^{-2}\log n)$ traces are sufficient for random graphs drawn from ${\mathcal{G}}(n,1/2)$ (i.e., $n$-vertex graphs where edges are present independently with probability $1/2$), assuming $p_{v}=\Omega(n^{-1/6}\log^{2/3}n)$. They also showed that $2^{\Omega(n)}$ traces are necessary to distinguish arbitrary graphs even when $p_{v}=1/2$.
The graph reconstruction problem outlined above is partially inspired by the analogous problem for binary strings, initially proposed by Batu et al. [2] and subsequently studied extensively [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. In the case of strings, the traces correspond to random subsequences (potentially subject to further noise). Despite extensive research, there is still a considerable gap between the best known upper and lower bounds on the sample complexity, whether the unknown string is arbitrary or random. The other motivation for our problem is the graph reconstruction problem from classical structural graph theory. There, the objective is to reconstruct an undirected $n$-vertex graph from the multiset of its induced subgraphs on $(n-1)$ vertices. Determining whether this is possible for arbitrary graphs is a famous unsolved problem [17, 18].
@@ I-A Our Results
Then, in the edge deletion model, $4p_{v}^{-2}p_{e}^{-1}\log n$ traces are sufficient to reconstruct $G$ with probability at least $1-1/n$, where the probability is also taken over the random choice of $G$. In the edge flip model, the corresponding bound is $12p_{v}^{-2}(1/2-f_{e})^{-2}\log n$.
@@ II-1 Notation and Conventions
Let $[k]$ denote the set $\{1,\ldots,k\}$ and, for a set $S$, let $\binom{S}{k}$ denote all subsets with cardinality $k$. We only consider undirected graphs $G=(V,E)$. We use $(u,v)$ to denote an edge, and $\{u,v\}$ to denote a pair in $\binom{V}{2}$, regardless of whether $(u,v)\in E$ or not. For $v\in V$, $\Gamma_{G}(v)$ denotes the neighborhood $\{v^{\prime}\in V:(v,v^{\prime})\in E\}$.
@@ II-2 Correlated Bits and Concentration Bounds
We will use standard notations from probability and statistics, e.g., $X\sim\mathrm{Bin}(N,\gamma)$ will mean the random variable $X$ is distributed according to the binomial distribution with parameters $N$ and $\gamma$. The following lemma will be used throughout the paper to quantify the probability that given two traces containing vertices $u$ and $v$, the edge $(u,v)$ is present in exactly one of them.
@@ III Reconstructing Random Graphs:
Algorithm 1 describes our procedure for pairing two traces by matching their common vertices. Informally, given two traces $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$, we find two induced subgraphs $G_{1}[S]$ and $G_{2}[T]$ with $|S|=|T|=k$, that are as close to being isomorphic as possible; specifically, we match their vertices in a way that minimizes the number of vertex pairs that induce an edge in one but not in the other. We set $k$ sufficiently large such that $k\approx|V_{1}\cap V_{2}|$. Our analysis shows that this process is guaranteed to find a large subset of the intersection $V_{1}\cap V_{2}$. We then augment the bijection to also match the remaining vertices in $V_{1}\cap V_{2}$. To do this, for each $v\in V_{1}$ and $v^{\prime}\in V_{2}$, we generate a signature based on $S$ and $T$ respectively, and match $v$ and $v^{\prime}$ iff their signatures are sufficiently similar. The signature is a binary vector that encodes the neighbors and non-neighbors of $v$ (resp. $v^{\prime}$) amongst $S$ (resp. $T$). The intuition is that these vectors are sufficiently similar iff $v$ and $v^{\prime}$ correspond to the same vertex in $G$.
@@ III-A Analysis
Note that if $A_{1}$ occurs, then $|V_{1}|$ and $|V_{2}|$ both have size at least $|V_{1}\cap V_{2}|\geq k$, and so there exists at least one triple $(S,T,\pi)$ where $S\subset V_{1},T\subset V_{2}$, and $|S|=|T|=k$; in other words, step 2 of the algorithm returns some triple. Let $\mathcal{T}$ denote the set of such triples. We next argue that with high probability the triple $(S^{\ast},T^{\ast},\pi^{\ast})$ that minimizes $\Delta_{\pi^{\ast}}$ has mostly fixed points. To do this, we define a mapping on triples $f:\mathcal{T}\rightarrow\mathcal{T}$ where $f(S,T,\pi)=(S^{\prime},T^{\prime},\pi^{\prime})$, where $S^{\prime}$ is an arbitrary set of vertices satisfying:
@@
Let $A_{2}$ be the event that the triple in $\mathcal{T}$ minimizing $\Delta$ has no non-fixed points. Then $\Pr[A_{2}|A_{1}]\geq 1-4n\exp(-kp_{e}^{3}/2592+2r\log n)$ assuming $25924\log n\leq k\cdot p_{e}^{3}/$ 2592.
@@
By Lemma III.3, for any triple in $\mathcal{T}$ with at least $b$ non-fixed points, there exists another triple with all fixed points that has a smaller value of $\Delta$ with probability at least $1-4\exp(-kbp_{e}^{3}/1296)$. So the probability there are any non-fixed points is at most $\sum_{b=1}^{n}4\cdot\exp(-kbp_{e}^{3}/1296+4b\log n+2r\log n)$, by the union bound. If $25924\log n\leq kp_{e}^{3}/$ 2592 then this is at most $4n\exp(-kp_{e}^{3}/2592+2r\log n)$. ∎
@@
Note that the last two of these imply that $kp_{e}^{3}=\omega(\log n)$, and so for large enough $n$, $25924\log n\leq kp_{e}^{3}/$ 2592, so the conditional in Theorem III.1 applies. Therefore, using Lemma III.2 and Theorems III.1 and III.2, we have:
@@ IV Reconstructing Random Graphs:
    The lower bound on the probability of $A_{2}$ in Theorem III.1 becomes $1-4n\exp(-kf_{e}(1/2-f_{e})^{4}/96+2r\log n)$ assuming $964\log n\leq kf_{e}(1/2-f_{e})^{4}/$ 96.
@@
Modifying the pairing proceduring (Lemma III.1) is slightly more involved. We now need enough traces so that for each pair of vertices $\{u,v\}\in\binom{V}{2}$, the majority of traces containing both nodes contain the edge $(u,v)$ iff $(u,v)\in E$.
@@ V Lower Bounds for Arbitrary Graphs
Consider the graphs $C_{n}$, an $n$-cycle, and $C_{n/2}+C_{n/2}$, the disjoint union of two $(n/2)$-cycles. Then, $\exp(\Omega(n))$ traces are necessary to distinguish them with constant probability, in the edge deletion model with $p_{v}=1,p_{e}=1/2$.
@@ Appendix A Omitted Proofs from Section II
since after fixing $W_{2}/W_{1}=w\in\{-1,1\}$, $Z_{1}/Z_{2}$ is uniform in $\{-1,1\}$ and hence equals $w$ with probability $1/2$. Lastly,
@@
For $21/4\leq f_{e}\leq 1/$ 2 we have $0\leq\rho_{2}/\rho_{1}-1\leq 1$. By applying the Chernoff bound, we get:
@@ Appendix B Omitted Proofs from Section III
Each node in $G_{\mathcal{N}}$ contributes 00 or $1$ to the flipped pair count. Observe that if two nodes in $G_{\mathcal{N}}$ are not the endpoints of an arc, then their contributions to the flipped pair count are independent of each other, as they correspond to two distinct pairs in ${\mathcal{N}}$, which map to two other pairs (in $\binom{V}{2}$ but not necessarily in ${\mathcal{N}}$). Therefore, an independent set in $G_{\mathcal{N}}$ corresponds to pairs in ${\mathcal{N}}$ that contribute 00 or $1$ to $\Delta_{\pi}$ independently of each other. What is the probability of such a pair contributing $1$ to $\Delta_{\pi}$? Either the pair was an edge (probability $p_{e}/2$) that goes to a non-edge (probability $(1-p_{e}/2)$), or it was a non-edge (probability $(1-p_{e}/2)$) that goes to an edge (probability $p_{e}/2$), for a total probability of $p_{e}(1-p_{e}/2)=\gamma_{4}$. Therefore, the contribution from any independent set of size $\ell$ in $G_{\mathcal{N}}$ is distributed as $\mathrm{Bin}(\ell,\gamma_{4})$.
@@
We claim first that every node in $G_{\mathcal{N}}$ has in-degree at most $1$ and out-degree at most $1$. Otherwise, some node corresponds to a pair in ${\mathcal{N}}$ that appears in the domain or image of $\sigma_{\pi}$ twice, contradicting the fact that $\sigma_{\pi}$ is a bijection. Therefore, $G_{\mathcal{N}}$ is a disjoint union of directed paths and directed cycles (including, possibly, $2$-cycles).
@@ Appendix C Omitted Proofs from Section V
Suppose that the vertices of $G_{1}=C_{n}$ are numbered $u_{1},\ldots,u_{n}$ in order, and suppose the vertices of $G_{2}=C_{n/2}+C_{n/2}$ are labeled $v_{1},\ldots,v_{n/2}$ in one cycle, and $v^{\prime}_{1},\ldots,v^{\prime}_{n/2}$ in the other cycle. Let $D_{1}$ (resp. $D_{2}$) be the distribution over traces from $C_{n}$ (resp. $C_{n/2}+C_{n/2}$). Consider the following coupling argument applied to $G_{1}$ and $G_{2}$ simultaneously. For $i=1$ to $n/2$, set $d_{i}=1$ with probability $1/4$ and 0 otherwise and if $d_{i}=1$ delete the following four edges:
@@
For a subset of vertices $S$, define $\textup{cut}(S)$ to be the number of edges with exactly one endpoint in $S$. Define the $t$-cut distribution to be the multiset $\{\textup{cut}(S):|S|=t\}$. For $t=2$, let $c_{1},c_{2},\ldots,c_{\binom{n}{2}}$ be the entries of this multiset. Consider the random process of generating a trace (recall that we are assuming that $p_{v}=1$ and $p_{e}=1/2$), randomly selecting two vertices $S=\{u,v\}$ in the trace, and returning the number of edges in the trace that have exactly one endpoint in $S$. Call this number $\tilde{c}(S)$ and note that $\tilde{c}(S)\sim\mathrm{Bin}(c(S),1/2)$. Since $S$ is equally likely to be any of the $\binom{n}{2}$ pairs of vertices, this process is equivalent to drawing a value $c_{i}$ from the multi-set $\{c_{1},c_{2},\ldots,c_{\binom{n}{2}}\}$, and then return a value drawn from $\mathrm{Bin}(c_{i},1/2)$. As such, the random output is a draw from a mixture of binomials and we can use an algorithm by Krishnamurthy et al. [10] to learn the 2-cut distribution using $\exp(O(n^{1/3}\log^{2/3}))$ traces. Note that this would allow us to distinguish $G_{1}$ and $G_{2}$ since these graphs have different $2$-cut distributions. Specifically, the number of sets $S\subset\binom{V}{2}$ such that $c(S)=2(n-2)$ is $r$ in the case of $G_{1}$ and $r-1$ in the case of $G_{2}$. Hence, it is possible to distinguish $G_{1}$ and $G_{2}$ with $\exp(O(n^{1/3}\log^{2/3}n))$ traces. ∎

### v2

@@ I Introduction
We are interested in the number of independently generated traces that are necessary to reconstruct the graph $G$ (with high probability). We refer to this number as the sample complexity of reconstruction. The problem was studied by McGregor and Sengupta [1], who considered the noiseless setting where $p_{e}=1$ (or equivalently $f_{e}=0$). They showed that $O(p_{v}^{-2}\log n)$ traces are sufficient for random graphs drawn from ${\mathcal{G}}(n,1/2)$ (i.e., $n$-vertex graphs where edges are present independently with probability 1/21/21 / 2), assuming $p_{v}=\Omega(n^{-1/6}\log^{2/3}n)$. They also showed that $2^{\Omega(n)}$ traces are necessary to distinguish arbitrary graphs even when $p_{v}=1/2$.
[…]
@@ I-A Our Results
[…]
@@ II-1 Notation and Conventions
[…]
@@ II-2 Correlated Bits and Concentration Bounds
[…]
@@ III Reconstructing Random Graphs:
[…]
@@ III-A Analysis
[…]
@@
Let $A_{2}$ be the event that the triple in $\mathcal{T}$ minimizing $\Delta$ has no non-fixed points. Then $\Pr[A_{2}|A_{1}]\geq 1-4n\exp(-kp_{e}^{3}/2592+2r\log n)$ assuming $4\log n\leq k\cdot p_{e}^{3}/2592$.
@@
By Lemma III.3, for any triple in $\mathcal{T}$ with at least $b$ non-fixed points, there exists another triple with all fixed points that has a smaller value of $\Delta$ with probability at least $1-4\exp(-kbp_{e}^{3}/1296)$. So the probability there are any non-fixed points is at most $\sum_{b=1}^{n}4\cdot\exp(-kbp_{e}^{3}/1296+4b\log n+2r\log n)$, by the union bound. If $4\log n\leq kp_{e}^{3}/2592$ then this is at most $4n\exp(-kp_{e}^{3}/2592+2r\log n)$. ∎
@@
Note that the last two of these imply that $kp_{e}^{3}=\omega(\log n)$, and so for large enough $n$, $4\log n\leq kp_{e}^{3}/2592$, so the conditional in Theorem III.1 applies. Therefore, using Lemma III.2 and Theorems III.1 and III.2, we have:
@@ IV Reconstructing Random Graphs:
    The lower bound on the probability of $A_{2}$ in Theorem III.1 becomes $1-4n\exp(-kf_{e}(1/2-f_{e})^{4}/96+2r\log n)$ assuming $4\log n\leq kf_{e}(1/2-f_{e})^{4}/96$.
@@
[…]
@@ V Lower Bounds for Arbitrary Graphs
[…]
@@ Appendix A Omitted Proofs from Section II
since after fixing $W_{2}/W_{1}=w\in\{-1,1\}$, $Z_{1}/Z_{2}$ is uniform in {$,1\}$ 1 } and hence equals $w$ with probability 1/21/21 / 2. Lastly,
@@
For $1/4\leq f_{e}\leq 1/2$ we have $0\leq\rho_{2}/\rho_{1}-1\leq 1$. By applying the Chernoff bound, we get:
@@ Appendix B Omitted Proofs from Section III
Each node in $G_{\mathcal{N}}$ contributes 0 or 111 to the flipped pair count. Observe that if two nodes in $G_{\mathcal{N}}$ are not the endpoints of an arc, then their contributions to the flipped pair count are independent of each other, as they correspond to two distinct pairs in ${\mathcal{N}}$, which map to two other pairs (in $\binom{V}{2}$ but not necessarily in ${\mathcal{N}}$). Therefore, an independent set in $G_{\mathcal{N}}$ corresponds to pairs in ${\mathcal{N}}$ that contribute 0 or 111 to $\Delta_{\pi}$ independently of each other. What is the probability of such a pair contributing 111 to $\Delta_{\pi}$? Either the pair was an edge (probability $p_{e}/2$) that goes to a non-edge (probability $(1-p_{e}/2)$), or it was a non-edge (probability $(1-p_{e}/2)$) that goes to an edge (probability $p_{e}/2$), for a total probability of $p_{e}(1-p_{e}/2)=\gamma_{4}$. Therefore, the contribution from any independent set of size $\ell$ in $G_{\mathcal{N}}$ is distributed as $\mathrm{Bin}(\ell,\gamma_{4})$.
@@
We claim first that every node in $G_{\mathcal{N}}$ has in-degree at most 111 and out-degree at most 111. Otherwise, some node corresponds to a pair in ${\mathcal{N}}$ that appears in the domain or image of $\sigma_{\pi}$ twice, contradicting the fact that $\sigma_{\pi}$ is a bijection. Therefore, $G_{\mathcal{N}}$ is a disjoint union of directed paths and directed cycles (including, possibly, 222-cycles).
@@ Appendix C Omitted Proofs from Section V
Suppose that the vertices of $G_{1}=C_{n}$ are numbered $u_{1},\ldots,u_{n}$ in order, and suppose the vertices of $G_{2}=C_{n/2}+C_{n/2}$ are labeled $v_{1},\ldots,v_{n/2}$ in one cycle, and $v^{\prime}_{1},\ldots,v^{\prime}_{n/2}$ in the other cycle. Let $D_{1}$ (resp. $D_{2}$) be the distribution over traces from $C_{n}$ (resp. $C_{n/2}+C_{n/2}$). Consider the following coupling argument applied to $G_{1}$ and $G_{2}$ simultaneously. For $i=1$ to $n/2$, set $d_{i}=1$ with probability 1/41/41 / 4 and 0 otherwise and if $d_{i}=1$ delete the following four edges:
@@
For a subset of vertices $S$, define $\textup{cut}(S)$ to be the number of edges with exactly one endpoint in $S$. Define the $t$-cut distribution to be the multiset $\{\textup{cut}(S):|S|=t\}$. For $t=2$, let $c_{1},c_{2},\ldots,c_{\binom{n}{2}}$ be the entries of this multiset. Consider the random process of generating a trace (recall that we are assuming that $p_{v}=1$ and $p_{e}=1/2$), randomly selecting two vertices $S=\{u,v\}$ in the trace, and returning the number of edges in the trace that have exactly one endpoint in $S$. Call this number $\tilde{c}(S)$ and note that $\tilde{c}(S)\sim\mathrm{Bin}(c(S),1/2)$. Since $S$ is equally likely to be any of the $\binom{n}{2}$ pairs of vertices, this process is equivalent to drawing a value $c_{i}$ from the multi-set $\{c_{1},c_{2},\ldots,c_{\binom{n}{2}}\}$, and then return a value drawn from $\mathrm{Bin}(c_{i},1/2)$. As such, the random output is a draw from a mixture of binomials and we can use an algorithm by Krishnamurthy et al. [10] to learn the 2-cut distribution using $\exp(O(n^{1/3}\log^{2/3}))$ traces. Note that this would allow us to distinguish $G_{1}$ and $G_{2}$ since these graphs have different 222-cut distributions. Specifically, the number of sets $S\subset\binom{V}{2}$ such that $c(S)=2(n-2)$ is $r$ in the case of $G_{1}$ and $r-1$ in the case of $G_{2}$. Hence, it is possible to distinguish $G_{1}$ and $G_{2}$ with $\exp(O(n^{1/3}\log^{2/3}n))$ traces. ∎